# HelpSteer2-Preference: Complementing Ratings with Preferencesthanks: 1NVIDIA, 2Georgia Tech, work done during internship at NVIDIA

### v1

@@ Abstract
Reward models are critical for aligning models to follow instructions, and are typically trained following one of two popular paradigms: Bradley-Terry style or Regression style. However, there is a lack of evidence that either approach is better than the other, when adequately matched for data. This is primarily because these approaches require data collected in different (but incompatible) formats, meaning that adequately matched data is not available in existing public datasets. To tackle this problem, we release preference annotations (designed for Bradley-Terry training) to complement existing ratings (designed for Regression style training) in the HelpSteer2 dataset. To improve data interpretability, preference annotations are accompanied with human-written justifications. Using this data, we conduct the first head-to-head comparison of Bradley-Terry and Regression models when adequately matched for data. Based on insights derived from such a comparison, we propose a novel approach to combine Bradley-Terry and Regression reward modeling. A Llama-3.1-70B-Instruct model tuned with this approach scores 94.1 on RewardBench, emerging top of more than 140 reward models as of 1 Oct 2024. We also demonstrate the effectiveness of this reward model at aligning models to follow instructions in RLHF. We open-source this dataset (CC-BY-4.0 license) at https://huggingface.co/datasets/nvidia/HelpSteer2 and openly release the trained Reward Model at https://huggingface.co/nvidia/Llama-3.1-Nemotron-70B-Reward.
@@ 1 Introduction
While training a reward model that can accurately separate good responses from bad is a consensus goal, there is less agreement on the best path to get there. On one side are the traditional Bradley-Terry Style Reward Models first introduced by Bai et al. (2022) and Ouyang et al. (2022) which seek to maximize the gap in reward between chosen and rejected responses to the same prompt. On the other side are Regression style Reward Models introduced by Wang et al. (2023) and Wang et al. (2024a) and have lately been used to train some of the top models on RewardBench (AllenAI, 2024) such as ArmoRM (Wang et al., 2024b) and Nemotron-4-340B-Reward (Wang et al., 2024d). Regression Reward Models train the model to predict the score (often Likert-5) for a response to a particular prompt. For many researchers and practitioners, deciding which style of Reward Models to adopt is challenging due to the absence of empirical studies that directly compare them when appropriately matched for data, which means:
@@
3. High Quality Following the adage ‘Garbage in Garbage out’, the role that the data format (regression-style vs. preferences) plays may not reliably show when the quality of data collection is the bottleneck, as the signal-to-noise ratio will be too low for the model to learn anything useful. Shen et al. (2024) found that the Llama-2-7B-Chat model performs essentially at a random level (i.e. <55% for a binary choice task) on RewardBench Chat-Hard category (Lambert et al., 2024) when trained on Ultrafeedback (Cui et al., 2023) or HH-RLHF (Bai et al., 2022), which is attributed to noise in the ground truth labels. Similarly, Wang et al. (2024d) found that Llama-3-70B models trained on HH-RLHF (Bai et al., 2022), Open Assistant (Köpf et al., 2023) or HelpSteer (Wang et al., 2023) do not surpass 60% on the same evaluation. In contrast, Llama-3-70B trained on HelpSteer2 (Wang et al., 2024d) performs above 80%. Without meeting a high bar on data quality, it may not be possible to discern the advantages of a particular data annotation methodology over another.
@@
3. From insights derived from the above comparison, we propose a novel approach to combine Bradley-Terry and Regression Reward Models, which can be used to train a reward model that scores 94.1 on RewardBench, which is the best performing model as of 1 Oct 2024. We also show how this reward model can be useful for RLHF.
@@ 2 Dataset
The last option (“Neither response is valid”) is to be used when both responses are so bad that trying to identify a winner is pointless. Inspired by Bai et al. (2022) and Touvron et al. (2023), we force the annotator to make a preference choice between the two responses (i.e. no option for Both responses are equally good), except for “Neither response is valid”. Our motivation was to reduce sitting-on-the-fence behavior and encourage annotators to look closer at differences (even if they appear to be minor), in order to much more robust preference information. The preference ranking guidelines and examples provided to annotators are in the Appendix B.
@@
Figure 1: Distribution of preferences between responses in HelpSteer2Preference against the difference in helpfulness scores between them from HelpSteer2. For clarity, A refers to Response 1 while B refers to Response 2. >>>>>>> > > means much better, >>>>> > means better, $>$ means slightly better and $=$ means as good as.
@@ 3.2 Training
Given that HelpSteer2-Preference contains not only the direction of preference between two responses but the magnitude ($m$) of this preference (1 - slightly better, 2 - better, 3 - much better), we also experiment with a Bradley-Terry with Margin loss introduced by (Touvron et al., 2023) in Eq. 2, thereafter referred to as Margin BT.
@@
We also introduce a new loss function named Scaled Bradley-Terry in Eq. 3, hereafter referred to as Scaled BT. Similar to Margin BT, its motivation lies in utilizing the preference magnitude information. However, we use the margin term outside of the log-sigmoid function rather than inside. From the perspective of data utilization, this can be viewed as a repeated sampling of response pairs for which the preference magnitude is higher. From the perspective of model training, this can be seen as requiring models to learn more (i.e. larger update) from responses that we know to be greatly different from each other. Unlike Margin BT, Scaled BT does not assume that the distance between the chosen and rejected rewards needs to be as least as large as the margin term.
@@ 4 Reward Model Results
While SteerLM Regression and Bradley-Terry models are separately comparable, we show that they have a synergistic effect and result in a stronger reward model when combined. Specifically, when initialized with the Helpfulness-only Regression model, a Scaled Bradley-Terry can reach overall RewardBench of 93.7. This is likely a result of the HelpSteer2 and HelpSteer2-Preference datasets containing complementary information, as first indicated by Sec. 2. Conceptually, this synergy is similar to the two-stage approach in performing preference tuning (i.e. Proximal Policy Optimization or Direct Preference Optimization) after doing supervised-finetuning. In addition, we found ExPO (Zheng et al., 2024) to be a simple and effective way of extrapolating the delta weights to further improve the model. Specifically, we use the Helpfulness-Only SteerLM Regression model as the weak model and the Scaled BT model (initialized with the Helpfulness-Only SteerLM Regression model) as the strong model. We then did an extrapolation factor search between 1.1 and 2.0 at intervals of 0.1. Upon finding 1.6 to be optimal, we searched between 1.51 and 1.69 at intervals of 0.01. The final extrapolation factor was 1.52. Neither the Regular BT model nor the Margin BT model improved upon the Helpfulness-only Regression Model that they were initialized with and therefore we did not try ExPO on them.
@@
Compared to SteerLM Regression and Bradley-Terry models that can score each sample independently, Pairwise Justifier models which involve choosing a better response between two candidates generally perform worse, with an overall RewardBench score of 90.0 or lower. We suspect that this is because this task formulation is much harder (implicitly involving scoring Response 1, then scoring Response 2 and finally determining which score is higher). This is also supported by the observation that strong external baseline models using this Pairwise Justifier approach (e.g. gpt-4o-2024-08-06 and Meta-Llama-3.1-70B-Instruct) score 86.7 and 84.0, which are similar to our Pairwise Justifier models. On the other hand, SteerLM Regression and Bradley-Terry models both decompose this problem into much simpler objectives that can be directly optimized towards.
@@ 5.3 Results
Across three DPO variants, there is consistent improvement over the base Llama-3.1-Instruct model in terms of GPT-4-Turbo MT-Bench as well as AlpacaEval 2.0 LC. We find that Scaled DPO performs best, underscoring the importance of adequately modelling the preference strength information we collected. However, no version of DPO can beat PPO or REINFORCE (across any of the three alignment metrics), highlighting the importance of using online training along with reward information.
@@
Our best model (trained with REINFORCE) achieves competitive performance with frontier models such as GPT-4o and Claude 3.5 Sonnet across the three popular alignment benchmarks (MT Bench, AlpacaEval 2.0 LC and Arena Hard). This highlights the strength of our Reward Model to guide online RLHF. REINFORCE significantly increases the response length compared to the base models, but its high length-controlled AlpacaEval 2.0 score suggests that the extra tokens are meaningful. However, it is important to bear in mind the limitations of these automated evaluations, as all of them use a LLM-as-a-judge (GPT-4-Turbo), which is generally perceived as weaker than the top-performing External Baselines like GPT-4o and Claude 3.5 Sonnet. As a result, the LLM-judge might not be able to adequately evaluate the quality of such model responses, in relation to responses generated by our aligned models. This means that such automated evaluation for alignment may be insufficient to inform on how our models perform in relation to such frontier models in diverse use cases. Given that our main purpose is to demonstrate how our trained reward models can be useful for RLHF, we leave more holistic evaluations as future work.
@@
6 Conclusion
We discover that Bradley-Terry style and Regression style Reward Models are competitive to one another when the optimal formulation of each style of Reward Models is used (e.g. Scaled Bradley-Terry or Helpfulness-Only SteerLM Regression). Furthermore, they can complement one another and training a scaled Bradley-Terry model, initialized on a Helpfulness-Only SteerLM Regression model achieves 94.1 on RewardBench overall, which is #1 on RewardBench leaderboard as of 1 Oct 2024. Finally, this reward model proves useful for aligning models to follow instruction using Online RLHF, particularly with the REINFORCE algorithm.
@@ Appendix A Related Works
There has been a series of work from sociological statistics on comparing the relative merits of ranking and rating style questionnaires in understanding human judgements. Alwin & Krosnick (1985) found both ranking and rating can perform similarly in terms of modelling overall preference but capture different (but complementary) aspects of their preference. McCarty & Shrum (2000) similarly found that forced-choice ranking yielded more differentiated preferwnce data compared to Likert-type ratings, with the additional advantage of fewer extreme scores. We build upon this literature to examine whether the same differences can be observed in collecting domain-general data for preference tuning of LLMs.
@@
OpenAssistant (Köpf et al., 2023) is a prominent crowd-sourced, open-source general domain human-annotated dataset with >10,000 conversation trees released by the Open Assistant organization. Open Assistant contains Likert-5 annotations for various attributes including helpfulness, creativity, and humor. Ultrafeedback (Cui et al., 2023) is a GPT-4 annotated dataset containing 64,000 prompts each with 4 responses. For each response, there are 4 attributes annotated on a Likert-10 scale: helpfulness, honesty, instruction-following, and truthfulness. With each rating, there is also a short (2-3 sentences) rationale that explains the rating. Ultrafeedback has previously been converted into a preference dataset (Tunstall et al., 2023) based on the mean of the 4 attribute scores - with the highest scoring response as the chosen and one of the other three responses selected at random as the rejected. HelpSteer (Wang et al., 2023) is a human-annotated general domain dataset with 10,000 prompts, each with 4 responses. Each response is labelled by 1 annotator on five Likert-5 attributes - helpfulness, correctness, coherence, complexity, and verbosity. However, Shen et al. (2024); Wang et al. (2024d) found the above datasets to substantially noisy, which means that reward models are unable to different between responses are not drastically different in quality. HelpSteer2 (Wang et al., 2024d) is a follow-up to HelpSteer with 10,000 real-world prompts, each with 2 responses. Each response is annotated on the same five Likert-5 attributes. However, every response is annotated by 3-5 annotators with the final ratings obtained by removing outlier annotations and then taking the average of the remaining annotations.
@@ Appendix E Training Hyper-parameters
[…]
@@
[…]
@@ Appendix H Aligned Model Evaluation Details
Arena Hard We follow (Dong et al., 2024; Meng et al., 2024; Wang et al., 2024d) to use Arena Hard (Li et al., 2024). Arena Hard contains 500 first-turn instructions obtained from challenging user queries on Chat Arena (Chiang et al., 2024). Prompts are classified using an LLM (Llama-3-70B-Instruct) to determine if they are complex, specific, real-world-application-related or require domain knowledge, problem-solving, creativity, technical accuracy. Prompts that meet at least 6 out of 7 criteria are labelled as challenging. Therefore, a huge proportion of prompts (>50%) are related to solving coding problems. Model responses are then compared with responses from GPT-4-0314 using GPT-4-1106-preview judge to calculate a win-rate over GPT-4-0314.

### v2

@@ Abstract
Reward models are critical for aligning models to follow instructions, and are typically trained following one of two popular paradigms: Bradley-Terry style or Regression style. However, there is a lack of evidence that either approach is better than the other, when adequately matched for data. This is primarily because these approaches require data collected in different (but incompatible) formats, meaning that adequately matched data is not available in existing public datasets. To tackle this problem, we release preference annotations (designed for Bradley-Terry training) to complement existing ratings (designed for Regression style training) in the HelpSteer2 dataset. To improve data interpretability, preference annotations are accompanied with human-written justifications. Using this data, we conduct the first head-to-head comparison of Bradley-Terry and Regression models when adequately matched for data. Based on insights derived from such a comparison, we propose a novel approach to combine Bradley-Terry and Regression reward modeling. A Llama-3.1-70B-Instruct model tuned with this approach scores 94.1 on RewardBench, emerging top of more than 140 reward models as of 1 Oct 2024. This reward model can then be used with REINFORCE algorithm (RLHF) to align an Instruct model to reach 85.0 on Arena Hard, which is No. 1 as of 1 Oct 2024.
[Uncaptioned image] Dataset (CC-BY-4.0-License): huggingface.co/datasets/nvidia/HelpSteer2
[Uncaptioned image] Reward Model: huggingface.co/nvidia/Llama-3.1-Nemotron-70B-Reward
[Uncaptioned image] Instruct Model: huggingface.co/nvidia/Llama-3.1-Nemotron-70B-Instruct
@@ 1 Introduction
While training a reward model that can accurately separate good responses from bad ones is a consensus goal, there is less agreement on the best path to get there. On one side are the traditional Bradley-Terry Style Reward Models first introduced by Bai et al. (2022) and Ouyang et al. (2022) which seek to maximize the gap in reward between chosen and rejected responses to the same prompt. On the other side are Regression style Reward Models introduced by Wang et al. (2024e) and Wang et al. (2024a) and have lately been used to train some of the top models on RewardBench (AllenAI, 2024) such as ArmoRM (Wang et al., 2024b) and Nemotron-4-340B-Reward (Wang et al., 2024d). Regression Reward Models train the model to predict the score (often Likert-5) for a response to a particular prompt. For many researchers and practitioners, deciding which style of Reward Models to adopt is challenging due to the absence of empirical studies that directly compare them when appropriately matched for data, which means:
@@
3. High Quality Following the adage ‘Garbage in Garbage out’, the role that the data format (regression-style vs. preferences) plays may not reliably show when the quality of data collection is the bottleneck, as the signal-to-noise ratio will be too low for the model to learn anything useful. Shen et al. (2024) found that the Llama-2-7B-Chat model performs essentially at a random level (i.e. <55% for a binary choice task) on RewardBench Chat-Hard category (Lambert et al., 2024) when trained on Ultrafeedback (Cui et al., 2023) or HH-RLHF (Bai et al., 2022), which is attributed to noise in the ground truth labels. Similarly, Wang et al. (2024d) found that Llama-3-70B models trained on HH-RLHF (Bai et al., 2022), Open Assistant (Köpf et al., 2023) or HelpSteer (Wang et al., 2024e) do not surpass 60% on the same evaluation. In contrast, Llama-3-70B trained on HelpSteer2 (Wang et al., 2024d) performs above 80%. Without meeting a high bar on data quality, it may not be possible to discern the advantages of a particular data annotation methodology over another.
@@
3. From insights derived from the above comparison, we propose a novel approach to combine Bradley-Terry and Regression Reward Models, which can be used to train a reward model that scores 94.1 on RewardBench, which is the best performing model as of 1 Oct 2024. This reward model can then be used with the REINFORCE algorithm (RLHF) to align a model to reach 85.0 on Arena Hard.
@@ 2 Dataset
The last option (“Neither response is valid”) is to be used when both responses are so bad that trying to identify a winner is pointless. Inspired by Bai et al. (2022) and Touvron et al. (2023), we force the annotator to make a preference choice between the two responses (i.e. no option for Both responses are equally good), except for “Neither response is valid”. Our motivation was to reduce sitting-on-the-fence behavior and encourage annotators to look closer at differences (even if they appear to be minor), in order to provide more robust preference information. The preference ranking guidelines and examples provided to annotators are in the Appendix B.
@@
Figure 1: Distribution of preferences between responses in HelpSteer2-Preference against the difference in helpfulness scores between them from HelpSteer2. For clarity, A refers to Response 1 while B refers to Response 2. $>$much-greater-thanabsent>>>> > > means much better, >>much-greater-than$>$ > means better, $>$ means slightly better and $=$ means as good as.
@@ 3.2 Training
Given that HelpSteer2-Preference contains not only the direction of preference between two responses but the magnitude ($m$) of this preference (1 - slightly better, 2 - better, 3 - much better), we also experiment with a Bradley-Terry with Margin loss introduced by Touvron et al. (2023) in Eq. 2, thereafter referred to as Margin BT.
@@
We also introduce a new loss function named Scaled Bradley-Terry in Eq. 3, hereafter referred to as Scaled BT. Similar to Margin BT, its motivation lies in utilizing the preference magnitude information. However, we use the margin term outside of the log-sigmoid function rather than inside. From the perspective of data utilization, this can be viewed as a repeated sampling of response pairs for which the preference magnitude is higher. From the perspective of model training, this can be seen as requiring models to learn more (i.e. larger update) from responses that we know to be greatly different from each other (more in Appendix J). Unlike Margin BT, Scaled BT does not assume that the distance between the chosen and rejected rewards needs to be as least as large as the margin term.
@@ 4 Reward Model Results
While SteerLM Regression and Bradley-Terry models are separately comparable, we show that they have a synergistic effect and result in a stronger reward model when combined. Specifically, when initialized with the Helpfulness-only Regression model, a Scaled Bradley-Terry can reach overall RewardBench of 93.7. This is likely a result of the HelpSteer2 and HelpSteer2-Preference datasets containing complementary information, as first indicated by Sec. 2. Conceptually, this synergy is similar to the two-stage approach in performing preference tuning (i.e. Proximal Policy Optimization or Direct Preference Optimization) after doing supervised-finetuning. In addition, we found ExPO (Zheng et al., 2024) to be a simple and effective way of extrapolating the delta weights to further improve the model. Specifically, we use the Helpfulness-Only SteerLM Regression model as the weak model and the Scaled BT model (initialized with the Helpfulness-Only SteerLM Regression model) as the strong model. We then did an extrapolation factor search between 1.1 and 2.0 at intervals of 0.1. Upon finding 1.6 to be optimal, we searched between 1.51 and 1.69 at intervals of 0.01. The final extrapolation factor was 1.52. Neither the Regular BT model nor the Margin BT model improved upon the Helpfulness-only Regression Model that they were initialized with and therefore we did not try ExPO on them. To better appreciate the contribution of certain design choices, we report ablation studies in Appendix K.
@@
Compared to SteerLM Regression and Bradley-Terry models that can score each sample independently, Pairwise Justifier models which involve choosing a better response between two candidates generally perform worse, with an overall RewardBench score of 90.0 or lower. We suspect that this is because this task formulation is much harder (implicitly involving scoring Response 1, then scoring Response 2 and finally determining which score is higher). This is also supported by the observation that strong external baseline models using this Pairwise Justifier approach (e.g. gpt-4o-2024-08-06 and Meta-Llama-3.1-70B-Instruct) score 86.7 and 84.0, which are similar to our Pairwise Justifier models. On the other hand, SteerLM Regression and Bradley-Terry models both decompose this problem into simpler objectives that can be directly optimized towards.
@@ 5.3 Results
Across three DPO variants, there is consistent improvement over the base Llama-3.1-Instruct model in terms of GPT-4-Turbo MT-Bench as well as AlpacaEval 2.0 LC. We find that Scaled DPO performs best, underscoring the importance of adequately modelling the preference strength information we collected. However, no version of DPO can beat PPO or REINFORCE (across any of the three alignment metrics), highlighting the importance of using online training along with reward information. We report ablation studies using alternative reward model in Appendix L.
@@
Our best model (trained with REINFORCE) achieves competitive performance with frontier models such as GPT-4o and Claude 3.5 Sonnet across the three popular alignment benchmarks (MT Bench, AlpacaEval 2.0 LC and Arena Hard). This highlights the strength of our Reward Model to guide online RLHF. REINFORCE significantly increases the response length compared to the base models, but its high length-controlled AlpacaEval 2.0 score suggests that the extra tokens are meaningful. However, it is important to bear in mind the limitations of these automated evaluations, as all of them use a LLM-as-a-judge (GPT-4-Turbo), which is generally perceived as weaker than the top-performing External Baselines like GPT-4o and Claude 3.5 Sonnet. As a result, the LLM-judge might not be able to adequately evaluate the quality of such model responses, in relation to responses generated by our aligned models. This means that such automated evaluation for alignment may be insufficient to inform on how our models perform in relation to such frontier models in diverse use cases. Given that our main purpose is to demonstrate how our trained reward models can be useful for RLHF, we leave more holistic evaluations to Appendix M.
@@ Appendix A Related Works
There has been a series of work from sociological statistics on comparing the relative merits of ranking and rating style questionnaires in understanding human judgements. Alwin & Krosnick (1985) found both ranking and rating can perform similarly in terms of modelling overall preference but capture different (but complementary) aspects of their preference. McCarty & Shrum (2000) similarly found that forced-choice ranking yielded more differentiated preference data compared to Likert-type ratings, with the additional advantage of fewer extreme scores. We build upon this literature to examine whether the same differences can be observed in collecting domain-general data for preference tuning of LLMs.
@@
OpenAssistant (Köpf et al., 2023) is a prominent crowd-sourced, open-source general domain human-annotated dataset with >10,000 conversation trees released by the Open Assistant organization. Open Assistant contains Likert-5 annotations for various attributes including helpfulness, creativity, and humor. Ultrafeedback (Cui et al., 2023) is a GPT-4 annotated dataset containing 64,000 prompts each with 4 responses. For each response, there are 4 attributes annotated on a Likert-10 scale: helpfulness, honesty, instruction-following, and truthfulness. With each rating, there is also a short (2-3 sentences) rationale that explains the rating. Ultrafeedback has previously been converted into a preference dataset (Tunstall et al., 2023) based on the mean of the 4 attribute scores - with the highest scoring response as the chosen and one of the other three responses selected at random as the rejected. HelpSteer (Wang et al., 2024e) is a human-annotated general domain dataset with 10,000 prompts, each with 4 responses. Each response is labelled by 1 annotator on five Likert-5 attributes - helpfulness, correctness, coherence, complexity, and verbosity. However, Shen et al. (2024); Wang et al. (2024d) found the above datasets to substantially noisy, which means that reward models are unable to different between responses are not drastically different in quality. HelpSteer2 (Wang et al., 2024d) is a follow-up to HelpSteer with 10,000 real-world prompts, each with 2 responses. Each response is annotated on the same five Likert-5 attributes. However, every response is annotated by 3-5 annotators with the final ratings obtained by removing outlier annotations and then taking the average of the remaining annotations.
@@ Appendix E Training Hyper-parameters
For SteerLM Regression models, we used 2 epochs of HelpSteer2 data, following Wang et al. (2024d). For Bradley-Terry models, we used 1 epoch of HelpSteer2-Preference data, as more than 1 epoch resulting in overfitting - similar to as observed by Zhu et al. (2024) and drastically higher validation loss and poorer RewardBench performance. For Pairwise Justification models, we train for 1 epoch for each setup - we also tried 2 epochs for initial experiments which showed minimal changes from only training on 1 epoch. For each experiment, we used global batch size of 128 using an AdamW optimizer with 10 warm-up steps and performed search over constant learning rates of {1, 2, 3}$e-6$. We report the optimal learning rate for each experiment in Table 1. We save checkpoints every 10 steps and evaluate three checkpoints with the lowest validation loss as well as the last checkpoint. Among them, we report the score with the highest RewardBench overall score. For SteerLM Regression Model with all five 5 HelpSteer attributes, we do grid search over all five attributes (between -1 and 1 at intervals of 0.1) over RewardBench.
@@
We trained DPO models for 2 epochs using the AdamW optimizer, after initially experimenting with 1 and 3 epochs but finding 2 epochs to be optimal. We use a global batch size of 64 response-pairs, weight decay of 0.01, with checkpoints saved every 10 steps, and 10 warmup steps followed by a constant learning rate. For hyper-parameter tuning, we performed a grid search over learning rates of {1, 2, 3}$e-7$ and KL penalties of {0.01, 0.001}. We report the optimal learning rate, KL penalty for each experiment in Table 4.
@@ Appendix H Aligned Model Evaluation Details
Arena Hard We follow Dong et al. (2024); Meng et al. (2024); Wang et al. (2024d) to use Arena Hard (Li et al., 2024). Arena Hard contains 500 first-turn instructions obtained from challenging user queries on Chat Arena (Chiang et al., 2024). Prompts are classified using an LLM (Llama-3-70B-Instruct) to determine if they are complex, specific, real-world-application-related or require domain knowledge, problem-solving, creativity, technical accuracy. Prompts that meet at least 6 out of 7 criteria are labelled as challenging. Therefore, a huge proportion of prompts (>50%) are related to solving coding problems. Model responses are then compared with responses from GPT-4-0314 using GPT-4-1106-preview judge to calculate a win-rate over GPT-4-0314.